# CodeWiki: Evaluating AI’s Ability to Generate Holistic Documentation for Large-Scale Codebases

### v4

@@ Abstract
Given a large and evolving codebase, the ability to automatically generate a holistic, architecture-aware documentation that captures not only individual functions but also their cross-file, cross-module, and system-level interactions remains an open challenge. Comprehensive documentation is essential for long-term software maintenance and team collaboration, yet current automated approaches still fail to model the rich semantic dependencies and architectural structures that define real-world software systems. We present CodeWiki, a unified framework for automated repository-level documentation across seven programming languages. CodeWiki introduces three key innovations: (i) hierarchical decomposition that preserves architectural context across multiple levels of granularity, (ii) recursive multi-agent processing with dynamic task delegation for scalable generation, and (iii) multi-modal synthesis that integrates textual descriptions with visual artifacts such as architecture diagrams and data-flow representations. To enable rigorous evaluation, we introduce CodeWikiBench, a comprehensive benchmark featuring multi-dimensional rubrics and LLM-based assessment protocols. Experimental results demonstrate that CodeWiki achieves a 68.79% quality score with proprietary models, outperforming the closed-source DeepWiki baseline (64.06%) by 4.73%, with particularly strong improvements on high-level scripting languages (+10.47%). We open-source CodeWiki to foster future research and community adoption.
[…]
@@ 1 Introduction
In the rapidly evolving landscape of software development, maintaining comprehensive and up-to-date documentation has become increasingly challenging as codebases grow in size and complexity. Recent industry surveys reveal that documentation represents a critical bottleneck in modern development workflows: approximately 31% of developers currently rely heavily on AI for documenting code, while 25% use AI extensively for creating and maintaining documentation—making documentation tasks among the highest priorities for AI integration over the next 3-5 years [43]. Notably, while developers express high satisfaction with AI tools for searching and learning, they indicate plans to substantially increase AI adoption for documentation and testing tasks, suggesting current tools remain insufficient for these critical needs. This widespread focus on AI-driven documentation reflects its dual burden: while essential for software maintenance and collaboration [9, 58, 5], manual creation and maintenance remain labor-intensive and often neglected [30]. The emergence of Large Language Models (LLMs) [54, 17] has opened opportunities for automating documentation generation. However, current approaches still face two fundamental limitations.
Limitations of Scalability at the Repository Level
Recent advances in code understanding and natural language generation have demonstrated promising results in function-level and file-level documentation tasks [12, 20, 37, 28, 26, 1, 7, 27, 53] However, existing approaches face significant challenges when scaling to repository-level documentation.
Unlike function-level documentation that focuses on individual components, repository-level documentation task is to capture architectural patterns, cross-module interactions, data flows, and system-wide design decisions [39, 46]. Furthermore, the hierarchical nature of software systems requires documentation that can serve multiple audiences, from high-level stakeholders seeking architectural overviews to developers requiring detailed implementation guidance [31].
@@
$\hookrightarrow$ To address these challenges, we present CodeWiki, an open-source framework specifically designed to generate holistic, structured repository-level documentation that includes diverse content types such as system architecture diagrams, data flow visualizations, and sequence diagrams while scaling effectively for projects of arbitrary size. CodeWiki introduces several key innovations: a hierarchical decomposition strategy inspired by dynamic programming principles that breaks complex repositories into manageable modules while preserving architectural coherence, a recursive agentic system with dynamic delegation capabilities that enables adaptive processing based on module complexity, and a comprehensive documentation synthesis approach that generates multiple forms of visual and textual documentation reflecting fine-grained codebase collaboration patterns.
Limitations of the Evaluation Landscape
Despite progress, systematic benchmarks remain lacking. Function-level and file-level work relies on general-purpose metrics (BLEU, ROUGE) that fail to capture documentation quality nuances [40, 11]. Repository-level evaluation is particularly challenging as documentation admits multiple valid structures sharing underlying architectural content, highlighting the need for robust assessment beyond surface-level similarity.
@@
$\hookrightarrow$ To fill this gap, we introduce CodeWikiBench, a comprehensive benchmark specifically designed for repository-level documentation. Instead of relying on surface-level similarity, we derive hierarchical rubrics from the official documentation maintained by project developers, enabling the evaluation of underlying repository properties. For robustness, CodeWikiBench further adopts an agentic assessment framework in which judge agents evaluate generated documentation against rubric requirements and aggregate results hierarchically, producing both quality scores and reliability measures.
Limitations of Multilingual Support
@@
Using CodeWikiBench, we demonstrate CodeWiki achieves 68.79% quality score with proprietary models, outperforming closed-source DeepWiki, while open-source alternatives achieve 64.80%, producing comprehensive documentation with architectural diagrams, usage patterns, and dependency visualizations (Appendix A).
In summary, the main contributions of this work are:
@@
    CodeWiki, an open-source semi-agentic framework for scalable repository-level documentation with proprietary and open-source LLM backends.
@@
    CodeWikiBench, a comprehensive repository-level documentation benchmark with hierarchical rubrics and agentic assessment.
@@
    Comprehensive multilingual repository documentation supporting 7 widely used languages.
@@
    Competitive performance on real-world repositories, improving over DeepWiki with both proprietary (Claude Sonnet 4) and open-source (Kimi K2 Instruct) models.
2 Related Work
2.1 Code Intelligence and LLMs
The rapid advancement of Large Language Models (LLMs) has revolutionized code intelligence tasks. Early work by von der Mosel et al. [47] demonstrated the effectiveness of Transformer-based models for software engineering, leading to specialized code models like CodeBERT [12], GraphCodeBERT [13], CodeT5 [49], CodeT5+ [50], and CodeTF [3]. These established early transformer-based approaches for code understanding and generation.
Comprehensive surveys by Niu et al. [34] and Hou et al. [19] established the importance of domain-specific pre-training, showing consistent improvements over general-purpose language models across code search, summarization, and generation tasks. However, recent studies highlight important limitations: Li et al. [22] found that models often make confident decisions without sufficient reasoning, while Zheng et al. [57] noted that LLM performance remains unstable and requires careful prompt engineering [42].
For code generation, HumanEval [6] and MBPP [2] established foundational benchmarks, later extended to multiple languages by MultiPL-E [4] and CodeScope [52]. More recently, CodeMMLU [32] has emerged as a comprehensive multi-task benchmark for assessing code understanding and reasoning capabilities of CodeLLMs.
2.2 Automated Documentation Generation
Documentation generation has emerged as a critical LLM application in software engineering. Traditional approaches focused on function-level summarization using datasets like CodeSearchNet [15] and FunCom [21]. While Sun et al. [45] found that ChatGPT can generate coherent comments, their quality often falls short of specialized models when evaluated using BLEU and ROUGE metrics.
Recent research has shifted toward repository-level documentation generation, which requires understanding complex dependencies and architectural context. Luo et al. [27] introduced RepoAgent, an LLM-powered framework that leverages static analysis and incremental processing. DocAgent [53] proposed a multi-agent system with specialized roles (Reader, Searcher, Writer, Verifier) for collaborative documentation generation. Both systems demonstrate significant improvements over baseline approaches but struggle with scaling to large repositories and maintaining cross-file consistency.
Evaluation methodologies have evolved from traditional NLP metrics like BLEU [35] and ROUGE [23] to more sophisticated approaches. Roy et al. [41] and Mastropaolo et al. [29] questioned the effectiveness of automatic metrics, while recent work has adopted LLM-as-a-judge methodologies [56, 25]. Multi-dimensional evaluation frameworks now assess completeness, helpfulness, and truthfulness, recognizing that documentation quality cannot be reduced to a single dimension.
2.3 Multi-Agent Systems for Software Engineering
Multi-agent architectures have gained prominence for tackling complex software engineering tasks requiring diverse expertise and collaborative problem-solving. AutoGen [51] introduced a framework enabling multi-agent communication with language models, allowing agents to assume different roles and collaborate on complex tasks. MetaGPT [18] demonstrated meta-programming for collaborative frameworks, where agents follow software development methodologies to produce comprehensive solutions.
Multi-agent approaches include HyperAgent [36], which presents generalist software engineering agents capable of solving coding tasks at scale, and AgileCoder [33], which introduces dynamic collaborative agents for software development based on agile methodology. In the code generation domain, MapCoder [55] employed map-reduce style collaboration to break down complex programming problems, while ChatDev [38] showed that multi-agent systems can manage the entire software development lifecycle. The integration of multi-agent systems with documentation represents a recent innovation, with DocAgent demonstrating how role-based collaboration can improve documentation quality through specialized agents for analysis, information retrieval, drafting, and verification. However, existing approaches often lack systematic dependency management and struggle with maintaining consistency across large codebases.
@@ 3.1 Repository Analysis and Hierarchical Module Decomposition
Following DocAgent [53], we employ Tree-Sitter parsers to extract ASTs, systematically identifying functions, methods, classes, structs, modules, and interdependencies: function calls, class inheritance, attribute access, and module imports. We normalize these to a unified depends_on relation for cross-language generalization, creating directed graph $G=(V,E)$ where edge $(u,v)\in E$ indicates component $u$ depends on $v$ [24].
@@
Topological sorting identifies zero-in-degree components—independent entry points where users interact (main functions, API endpoints, CLI, public interfaces). High-level components are hierarchically decomposed into manageable modules through recursive partitioning considering component interdependencies and semantic coherence. For scalability, only component IDs serve as input, resulting in a feature-oriented module tree.
@@ 3.2 Recursive Documentation Generation
CodeWiki’s core innovation is recursive agent-based processing, enabling arbitrary repository size handling while maintaining bounded complexity and architectural coherence.
@@
Each leaf module is assigned a specialized agent equipped with: (1) complete source code access, (2) full module tree for cross-module understanding, (3) documentation workspace tools (view, create, edit operations), and (4) dependency graph traversal for contextual exploration.
The agent workflow: analyzes components to understand functionality and interfaces, explores context by traversing dependencies within and across modules, and generates comprehensive markdown documentation including descriptions, usage examples, API specifications, and architectural insights.

### v5

@@ Abstract
Comprehensive software documentation is crucial yet costly to produce. Despite recent advances in large language models (LLMs), generating holistic, architecture-aware documentation at the repository level remains challenging due to complex and evolving codebases that exceed LLM context limits. Existing automated methods struggle to capture rich semantic dependencies and architectural structure. We present CodeWiki, a unified framework for automated repository-level documentation across seven mainstream programming languages. CodeWiki combines top-down hierarchical decomposition with a divide-and-conquer agent system to preserve architectural context and scale documentation generation, and a bottom-up synthesis that integrates textual descriptions with visual artifacts such as architecture and data-flow diagrams. We also introduce CodeWikiBench, a benchmark with hierarchical rubrics and LLM-based evaluation protocols. Experiments show that CodeWiki achieves a 68.79% quality score with proprietary models, outperforming the closed-source DeepWiki baseline by 4.73%, with especially strong gains on scripting languages. CodeWiki is released as open source to support future research.
@@ 1 Introduction
In the rapidly evolving landscape of software development, maintaining comprehensive and up-to-date documentation has become increasingly challenging as codebases grow in size and complexity. Studies indicate that developers spend approximately 58% of their working time on program comprehension activities (Xia et al., 2018), highlighting the critical importance of accessible and accurate documentation. Despite documentation being recognized as essential for software maintenance and collaboration (de Souza et al., 2005; Zhi et al., 2015; Chen and Huang, 2009), the manual creation and maintenance of repository-level documentation remain a labor-intensive and often neglected aspect of software engineering practice (McBurney et al., 2018). The emergence of Large Language Models (LLMs)  (Vaswani et al., 2017; Zhang et al., 2023; He et al., 2025b) has opened opportunities for automating documentation generation tasks. However, current approaches still face two fundamental limitations.
Limitations of Scalability at the Repository Level.
Recent advances in code understanding and natural language generation have demonstrated promising results in function and file documentation tasks (Feng et al., 2020; Khan and Uddin, 2023; Poudel et al., 2024; Makharev and Ivanov, 2025; Lomshakov et al., 2024; Ahmed and Devanbu, 2023; Choi et al., 2023; Luo et al., 2024; Yang et al., 2025). However, existing approaches face significant challenges when scaling to repository documentation. Unlike function documentation that focuses on individual components, repository documentation must capture architectural patterns, interactions between modules, data flows, and design decisions that span the entire system (Rai et al., 2022; Treude et al., 2020). Furthermore, the complex hierarchical nature of software systems requires documentation that can serve multiple purposes, such as for stakeholders seeking architectural overviews or developers requiring detailed implementation guidance (Nassif and Robillard, 2025).
@@
$\hookrightarrow$ To address these challenges, we present CodeWiki, an open-source framework designed to generate holistic repository documentation that includes diverse content types such as system architecture diagrams, data flow visualizations, and sequence diagrams while scaling effectively for projects of arbitrary size. While prior work such as RepoAgent (Luo et al., 2024) aggregates function and class documentation into repository structures, and open-source alternatives like Open-DeepWiki apply LLMs to entire repositories, CodeWiki differs through its hierarchical synthesis approach where parent modules are synthesized from child documentation, combined with integrated visual artifacts capturing architectural patterns. CodeWiki introduces three key innovations: (1) a hierarchical decomposition strategy inspired by dynamic programming principles that breaks complex repositories into manageable modules while preserving architectural coherence; (2) a scalable divide-and-conquer agent system that dynamically determines the relevant context for each module, recursively partitions it into sub-modules handled by dedicated sub-agents; and (3) a bottom-up synthesis process that integrates textual descriptions with visual artifacts such as architecture diagrams and data-flow representations to generate high-level documentation.
Limitations of the Evaluation Landscape.
Despite progress in automated documentation, there is still a lack of systematic and standardized benchmarks. Existing work at the function-level and file-level has mostly been evaluated using general-purpose text generation metrics (e.g., BLEU, ROUGE), which fail to capture the nuanced quality of software documentation (Roy et al., 2021; Evtikhiev et al., 2023). For repository-level documentation, the large scale and complexity of software systems make evaluation particularly challenging, as the quality of documentation cannot be adequately judged by surface-level text-similarity metrics alone. A single repository can be described through documentation with different structures and presentations, yet all share the same underlying architectural content. This highlights the lack of a robust evaluation system that can fairly assess documentation quality regardless of how it is presented.
@@
$\hookrightarrow$ To fill this gap, we introduce CodeWikiBench, the first benchmark specifically designed for repository-level documentation. Instead of relying on surface-level similarity, we automatically derive hierarchical evaluation rubrics from the ground-truth official documentation maintained by project developers, enabling the automated evaluation of machine-generated documentation using the derived rubrics. This approach is inspired by recent advances in hierarchical rubric-based evaluation (Starace et al., 2025), which demonstrates the effectiveness of decomposing complex evaluation tasks into smaller, well-defined sub-tasks with clear grading criteria. For robustness, CodeWikiBench adopts an agentic assessment framework in which judge agents evaluate generated documentation against rubric requirements and aggregate results hierarchically, producing both quality scores and reliability measures.
Limitations of Multilingual Support.
@@
Using CodeWikiBench, we evaluate the effectiveness of CodeWiki against multiple repository-level documentation systems, including open-source baselines (OpenDeepWiki [^1] https://github.com/AIDotNet/OpenDeepWiki, deepwiki-open [^2] https://github.com/AsyncFuncAI/deepwiki-open) and the closed-source DeepWiki. The results show that CodeWiki achieves an overall quality score of 68.79%, outperforming all baselines with improvements of up to 18.54% over the strongest baseline, DeepWiki, on individual repositories. CodeWiki produces comprehensive documentation encompassing architectural diagrams, usage patterns, and cross-module dependency visualizations (see Appendix A for an example).
In summary, our main contributions are:
@@
    CodeWiki, an open-source framework for scalable repository documentation generation combining top-down and bottom-up analyses.
@@
    CodeWikiBench, a benchmark for repository documentation featuring hierarchical rubrics and an agentic assessment framework.
@@
    Support for documentation generation across seven programming languages (Python, Java, JavaScript, TypeScript, C, C++, and C#).
@@
    Experimental validation demonstrating improvements over all baselines, with gains of up to 18.54% on individual repositories and 4.73% on average compared to the closed-source DeepWiki system.
A detailed discussion of related work and positioning relative to prior approaches is provided in Appendix B.
@@ 2.1 Repository Analysis and Hierarchical Module Decomposition
Following DocAgent (Yang et al., 2025), we employ Tree-Sitter parsers to extract Abstract Syntax Trees (ASTs). We systematically identify functions, methods, classes, structs, modules, and interdependencies including: function calls (runtime dependencies), class inheritance (structural hierarchies), attribute access (data dependencies), and module imports (compilation dependencies). We normalize these to a unified depends_on relation for cross-language generalization.
This creates a directed graph $G=(V,E)$ where vertices $V$ represent components and edge $(u,v)\in E$ indicates component $u$ depends on $v$. Understanding $v$ is prerequisite to comprehending $u$, fundamental to our hierarchical processing (Liu et al., 2023).
@@
Topological sorting identifies zero-in-degree components, independent entry points where users interact with the project (main functions, API endpoints, CLI, public interfaces). Components from these files represent high-level features and documentation starting points.
High-level components are hierarchically decomposed into manageable modules through recursive partitioning. For scalability, only component IDs (relative paths) serve as input. Decomposition considers: component interdependencies (grouping related functionality) and semantic coherence (preserving logical boundaries). The result is a feature-oriented module tree where each node contains component lists and sub-modules.
@@ 2.2 Recursive Documentation Generation
CodeWiki’s core innovation is recursive agent processing, enabling arbitrary repository size handling while maintaining bounded complexity and architectural coherence.
@@
Each leaf module is assigned a specialized agent equipped with: (1) complete source code access, (2) full module tree for cross-module understanding, (3) documentation workspace tools, and (4) dependency graph traversal for contextual exploration.
The agent workflow: analyzes components to understand functionality and interfaces, explores context by traversing dependencies, and generates comprehensive markdown documentation including descriptions, usage examples, API specifications, and architectural insights.